Jasmine and Jesse drive from Abingdon to Boulia at a speed of 90km/h, and then from Boulia to Clarenvale at an average speed of 100km/h. If the total time taken for the journey from Abingdon to Clarenvale was two hours and the average speed for the entire journey was 96km/h, how far is it from Abingdon to Boulia.
I have figured out from the given average speed of the entire journey and time that the entire journey was 196 km in distance, although I may be wrong. How do you work this out.

Answer by jorel555(1290)

If the average speed over two hours was 96 kmh, then the total distance was 2*96=192 kilometers. Let n be the time from Boulia to Clarenvale. Then:
100n+90(2-n)=192
10n+180=192
10n=12
n=1.2
2-n=.8
.8 * 90 kmh=72 kilometers. ☺☺☺☺